# PIF-Net: Ill-Posed Prior Guided Multispectral and Hyperspectral Image Fusion via Invertible Mamba and Fusion-Aware LoRA

### v1

@@ Introduction
Our model is implemented in PyTorch and trained on an NVIDIA A30 GPU. The hidden dimension $D$ of PIF-Net is set to 64, and the number of feature extraction blocks $L$ is 4. We use the $L_{1}$ loss function combined with the AdamW optimizer (Kingma and Ba 2015), starting with a learning rate of $1\times 10^{-4}$, which is halved every 200 epochs. The batch size is set to 8, and the total training process spans 500 epochs. This setup ensures stable convergence and efficient learning.
@@
To investigate the influence of the ill-posed residual prior on model performance, we introduce a weighting parameter $\beta$ to control the strength of this prior. A series of ablation experiments were conducted on the PaviaU ×4$\times$ 4× 4 dataset to evaluate the effect of varying $\beta$ values. The results, presented in Table 2, reveal a clear trend in model performance as $\beta$ changes. Specifically, the ill-posed residual prior significantly contributes to improving both model stability and fusion accuracy. The best performance is achieved when $\beta=1$. In contrast, setting $\beta$ to 0.4 or 0.8 results in no notable improvement and may even slightly degrade performance. Moreover, when $\beta=0$, indicating the model does not utilize the residual prior, performance is at its lowest, further underscoring the critical role of this prior.
Table 2: Impact of varying $\beta$ values on PSNR performance on the PaviaU ×4$\times$ 4× 4 dataset, demonstrating how the weighting of the ill-posed residual prior influences fusion quality. The best results are in bold.
@@
To assess the impact of the Invertible Mamba Block on fusion performance, we performed an ablation study on the PaviaU ×4$\times$ 4× 4 dataset, comparing the model’s performance with and without this block. As shown in Table 3, incorporating the Mamba Block leads to a notable improvement in fusion quality. Specifically, the PSNR increases from 37.1781 dB to 39.8246 dB, while SSIM improves from 0.9637 to 0.9845. These results highlight the effectiveness of the Mamba Block in enhancing feature representation and boosting the fidelity and quality of the fused images.
Table 3: Performance impact of the Invertible Mamba Block on the PaviaU ×4$\times$ 4× 4 dataset. The best results are in bold.
@@
We also investigate the influence of the FAM-LoRA module using a similar ablation study on the PaviaU ×4$\times$ 4× 4 dataset. As reported in Table 4, the inclusion of FAM-LoRA considerably enhances the model’s accuracy. Without this module, the PSNR and SSIM are 36.9211 dB and 0.9607, respectively. When FAM-LoRA is integrated, these metrics rise substantially to 39.8246 dB and 0.9845. This demonstrates that FAM-LoRA effectively facilitates feature aggregation, leading to superior fine-detail fusion and overall image quality.
Table 4: Performance impact of the FAM-LoRA module on the PaviaU ×4$\times$ 4× 4 dataset. The best results are in bold.
@@
Table 5: Ablation study on loss components ($\mathcal{L}_{1}$ loss always included) conducted on the PaviaU ×4$\times$ 4× 4 dataset. The best results are in bold.
@@
We comprehensively evaluate PIF-Net against several SOTA methods, with results in Table 1 showing its clear superiority across all metrics. PIF-Net consistently achieves the highest PSNR and SSIM and the lowest SAM and ERGAS across datasets and scales. Notably, on challenging ×4$\times$ 4× 4 and ×8$\times$ 8× 8 tasks, PIF-Net significantly outperforms competitors—for instance, on PaviaU ×4$\times$ 4× 4, it surpasses the second-best by 0.9045 dB PSNR, and on Chikusei ×8$\times$ 8× 8, it sets a new benchmark with 50.0124 dB PSNR. These results validate that the integration of the Invertible Mamba Block and FAM-LoRA effectively captures spectral-spatial correlations and preserves high-frequency details, empowering PIF-Net’s superior performance, especially at large scaling factors.
@@
To further evaluate visual quality, we qualitatively compare all methods on three datasets in Figure 5, showing pseudo-color images, ×4$\times$ 4× 4 super-resolution results, and SAM error maps. PIF-Net consistently delivers sharper textures, more coherent spectral structures, and fewer artifacts than state-of-the-art methods. Across Chikusei, PaviaU, and Houston scenes, it produces cleaner edges and better spectral preservation, evidenced by lower SAM errors. These qualitative results align with the quantitative gains in Table 1, confirming PIF-Net’s superior performance both visually and numerically.
Figure 6: Efficiency comparison of models on the PaviaU dataset for $\times$4 SR. Parameters (M) and inference time (ms, on NVIDIA A30) are shown in log scale. PSNR and SSIM indicate fusion quality. PIF-Net offers the best trade-off between accuracy and efficiency.
@@
We evaluated model efficiency on the PaviaU dataset for the ×4$\times$ 4× 4 super-resolution task, measuring inference time on an NVIDIA A30 GPU. Figure 6 shows parameters and inference time on a logarithmic scale. Benefiting from the novel architecture, invertible Mamba module, and lightweight multi-head LoRA, PIF-Net achieves an optimal trade-off among size, speed, and accuracy. With only 1.73 million parameters and 9.3 ms inference time, it attains the highest PSNR and SSIM, highlighting its efficiency and suitability for real-time use.
@@ Conclusion
This paper proposes PIF-Net, an innovative multimodal hyperspectral image fusion (MHIF) framework. Guided by ill-posed priors, the framework employs a dual-branch architecture that integrates Haar wavelet transform with an invertible Mamba module, achieving high-fidelity information transformation and stable transmission, while ensuring the invertibility of feature transformations and the stability of gradient flow. Meanwhile, by combining a spatial-domain fusion strategy based on low-frequency priors with a lightweight LoRA module, it effectively enhances the dynamic calibration of spatial details and spectral features. Furthermore, we innovatively design a guided feature consistency loss that significantly improves semantic alignment and representation across domains. Extensive experimental results demonstrate that PIF-Net achieves SOTA performance on three public datasets.

### v2

@@ Introduction
Our model is implemented in PyTorch and trained on an NVIDIA A30 GPU. The hidden dimension $D$ of PIF-Net is set to 64, and the number of feature extraction blocks $L$ is 4. We adopt the AdamW optimizer (adam) with an initial learning rate of $1\times 10^{-4}$, which is halved every 200 epochs. The batch size is 8, and the total training spans 500 epochs. For all comparison methods, losses follow their official default implementations.
@@
To investigate the influence of the ill-posed residual prior on model performance, we introduce a weighting parameter $\beta$ to control the strength of this prior. A series of ablation experiments were conducted on the PaviaU $\times 4$ dataset to evaluate the effect of varying $\beta$ values. The results, presented in Table 2, reveal a clear trend in model performance as $\beta$ changes. Specifically, the ill-posed residual prior significantly contributes to improving both model stability and fusion accuracy. The best performance is achieved when $\beta=1$. In contrast, setting $\beta$ to 0.4 or 0.8 results in no notable improvement and may even slightly degrade performance. Moreover, when $\beta=0$, indicating the model does not utilize the residual prior, performance is at its lowest, further underscoring the critical role of this prior.
@@
Table 2: Impact of varying $\beta$ values on PSNR performance on the PaviaU $\times 4$ dataset, demonstrating how the weighting of the ill-posed residual prior influences fusion quality. The best results are in bold.
@@
To assess the impact of the Invertible Mamba Block on fusion performance, we performed an ablation study on the PaviaU $\times 4$ dataset, comparing the model’s performance with and without this block. As shown in Table 3, incorporating the Mamba Block leads to a notable improvement in fusion quality. Specifically, the PSNR increases from 37.1781 dB to 39.8246 dB, while SSIM improves from 0.9637 to 0.9845. These results highlight the effectiveness of the Mamba Block in enhancing feature representation and boosting the fidelity and quality of the fused images.
@@
Table 3: Performance impact of the Invertible Mamba Block on the PaviaU $\times 4$ dataset. The best results are in bold.
@@
We also investigate the influence of the FAM-LoRA module using a similar ablation study on the PaviaU $\times 4$ dataset. As reported in Table 4, the inclusion of FAM-LoRA considerably enhances the model’s accuracy. Without this module, the PSNR and SSIM are 36.9211 dB and 0.9607, respectively. When FAM-LoRA is integrated, these metrics rise substantially to 39.8246 dB and 0.9845. This demonstrates that FAM-LoRA effectively facilitates feature aggregation, leading to superior fine-detail fusion and overall image quality.
@@
Table 4: Performance impact of the FAM-LoRA module on the PaviaU $\times 4$ dataset. The best results are in bold.
@@
Table 5: Ablation study on loss components ($\mathcal{L}_{1}$ loss always included) conducted on the PaviaU $\times 4$ dataset. The best results are in bold.
@@
We comprehensively evaluate PIF-Net against several SOTA methods, with results in Table 1 showing its clear superiority across all metrics. PIF-Net consistently achieves the highest PSNR and SSIM and the lowest SAM and ERGAS across datasets and scales. Notably, on challenging $\times 4$ and $\times 8$ tasks, PIF-Net significantly outperforms competitors—for instance, on PaviaU $\times 4$, it surpasses the second-best by 0.9045 dB PSNR, and on Chikusei $\times 8$, it sets a new benchmark with 50.0124 dB PSNR. These results validate that the integration of the Invertible Mamba Block and FAM-LoRA effectively captures spectral-spatial correlations and preserves high-frequency details, empowering PIF-Net’s superior performance, especially at large scaling factors
@@
We evaluate the computational efficiency of the models on the PaviaU dataset for the $\times$4 super-resolution task, measuring inference latency on an NVIDIA A30 GPU under identical implementation settings. As shown in Figure 6, both the number of parameters and the inference latency are presented on a logarithmic scale for clearer comparison across models. PIF-Net achieves a favorable trade-off among model size, inference speed, and reconstruction accuracy. With only 1.73 million parameters and a latency of 9.3 ms, PIF-Net achieves the highest PSNR and SSIM among all compared methods, demonstrating superior efficiency and strong potential for real-time applications.
Figure 6: Comparison of model efficiency on PaviaU for $\times 4$ super-resolution. Both the number of parameters (in millions) and inference latency (in milliseconds, measured on an NVIDIA A30 GPU) are displayed on logarithmic scales.
@@
We evaluate the computational efficiency of all competing methods on the PaviaU dataset for the $\times 4$ super-resolution task, measuring inference latency on an NVIDIA A30 GPU under identical implementation and batch settings. As illustrated in Figure 6, both model size (in millions of parameters) and inference latency (in milliseconds) are displayed on logarithmic scales to better highlight relative differences. PIF-Net achieves a compelling balance among model complexity, speed, and reconstruction fidelity. With only 1.73M parameters and a latency of 9.3 ms, it outperforms all other methods in terms of both PSNR and SSIM, demonstrating SOTA accuracy and high computational efficiency, which makes it well suited for practical and real-time fusion scenarios.
@@ Conclusion
This paper proposes PIF-Net, an innovative MHIF framework. Guided by ill-posed priors, the framework employs a dual-branch architecture that integrates HWT with Invertible Mamba Blocks, achieving high-fidelity information transformation and stable transmission, while ensuring the invertibility of feature transformations and the stability of gradient flow. Meanwhile, by combining a spatial-domain fusion strategy based on low-frequency priors with a lightweight LoRA fusion module, it effectively enhances the dynamic calibration of spatial details and spectral features. Furthermore, we innovatively design a guided feature consistency loss that improves semantic alignment and representation across domains. Extensive experimental results demonstrate that PIF-Net achieves SOTA performance on three public datasets.
Acknowledgments
This research was supported by the National Key Research and Development Program of China (No. 2023YFB4502304).